OpenStudy (anonymous):

Which of the following will give you the incorrect slope? 1) y2-y1 over x2 - x1 2) y2-y1 over x1 -x2 3) y1 - y2 over x1 - x2 4) rise over run

OpenStudy (anonymous):

the incorrect one which one is incorrect ???

OpenStudy (jhannybean):

I just said...

OpenStudy (jhannybean):

Or just take two points and input them into all the possible outcomes, and see which one does not match the original formula.

OpenStudy (jhannybean):

okay, let's take the point (2,3), (4,5)

OpenStudy (jhannybean):

And then input these two points into ALL your choices. all your choices should match the original slope except one.

OpenStudy (anonymous):

so what the incorrect one ???

OpenStudy (jhannybean):

let's go 1 by 1. use point (2,3) , (4,5) \[m=\frac{ y_{2}-y_{1} }{ x_{2}-x_{1} } = \frac{ 5-3 }{ 4-2 }=\frac{ 2 }{ 2 }=1\] so our slope should equal 1. 1) y2-y1/x2-x1 = 1, plugging in points for our original. RIGHT 2) y2-y1/x1-x2 = (5-3)/2-4 = 2/-2 = -1 INCORRECT 3) y1-y2/x1-x2 = (3-5)/(2-4) = (-2/-2) = 1 RIGHT 4) y1-y2/x1-x2 = (3-5)/(2-4) = (-2/-2) = 1 RIGHT

OpenStudy (anonymous):

ok thank you

OpenStudy (jhannybean):

Yep.